# Extrinsic contribution to bosonic thermal Hall transport

### v1

@@ Appendix E Solving Boltzmann’s equation and the displaced local equilibrium current
Here we derive the solution Eq.(34) to the equation Eq.(III.3), to the first order in $\hbar$, equivalently in coordinate shifts $\delta\mathsf{X}$. To this order the “compensation term” $\mathsf{g}_{\rm d}^{\rm sj}(X,p)$ and out-of-equilibrium part $\mathsf{g}_{\rm d}^{\rm rta}(X,p)$ can be solved for independently, which we do following the steps of Ref.Sinitsyn et al. (2005) adapted to our generalized bosonic problem and notations.
@@ E.1 Displaced local equilibrium current
×$\qquad\qquad\times\bigg{[}(\mathsf{g}_{\rm d})_{n,n}(X,p)-(\mathsf{g}_{\rm d})_{n^{\prime},n^{\prime}}(X,p^{\prime})-\delta\mathsf{X}^{\mu}_{n,n^{\prime}}(p,p^{\prime})\,\mathsf{\partial}_{X_{\mu}}T\;\partial_{T}f\!\left((\underline{\mathsf{K}}_{\rm d})_{n^{(}{}^{\prime}{}^{)},n^{(}{}^{\prime}{}^{)}}(X,p^{(}{}^{\prime}{}^{)})\right)\bigg{]}+\dots$
where “+$\dots$+ …” contains terms higher-order in powers of $\hbar$ or beyond linear response, and the optional set of apostrophes in $\partial_{T}f$ follows from Eq.(91).
We now use the fact that Eq.(26), up to the approximations discussed elsewhere, takes the form Eq.(34). Then, a solution to the collision integral is given by Eq.(35b) in the main text, which is clearly gauge invariant.
@@ E.2 Response in the relaxation time approximation (RTA)
×$\qquad\qquad\times\left|\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})\right|^{2}\times\bigg{[}(\mathsf{g}_{\rm d}^{\rm rta})_{n,n}(X,p)-(\mathsf{g}_{\rm d}^{\rm rta})_{n^{\prime},n^{\prime}}(X,p^{\prime})\bigg{]}.$
@@ F.1 Thermal Hall conductivity
Contrarily to the scalar fermionic case, because $\mathsf{\Omega}^{\mathsf{W}}_{p_{\mu}p_{\nu}}$ is not antisymmetric a priori, these two currents are not identical a priori. The local current due to disorder, ${J}_{y}^{\rm dis}\equiv{J}_{y}^{\rm sj,rta}+{J}_{y}^{\rm gv,sj}$, is thus
@@ F.2 Electrical Hall effect
Appendix G Applications – Details
G.1 Effective low-energy theory for magnons in an antiferromagnet
@@
In terms of $m_{i},n_{i}$ with $i=1,2$, and writing $\partial_{1}\equiv\partial_{x},\partial_{2}\equiv\partial_{y}$, the hamiltonian Eq.(V.1) reads
@@
G.2 Diagonalization of the clean problem
@@
Here we provide $\Psi_{1},\Psi_{2}$ to the first order in $\eta$ only. This is sufficient to compute the leading contribution to the Berry curvature $\mathsf{\Omega}_{p_{x}p_{y}}$ given above, Eq.(G.2), which is odd under time reversal, and thus odd in powers of $\eta$.
G.3 Explicit calculations in two given instances of disorder (b,c)
G.3.1 Case (b): a case with both intrinsic and extrinsic chirality
@@
[…]
@@
[…]
@@
[…]
@@
where #$\#_{\mu}(p)$ and #$\#^{\prime}_{\mu}(p)$ are real quantities which we do not write explicitly here.
G.3.2 Case (c): a case where the clean thermal Hall conductivity is zero
@@
$[\mathsf{S}^{-1}(p)\mathsf{W}(p)\mathsf{S}(p)]_{n,n}=0,$
@@
G.4 Concrete microscopic spin model: details
@@
where ${\mathtt{a}}$ is the lattice parameter and “+$\dots$+ …” contains terms that are higher order in gradients or beyond quadratic order in $\bm{m},\bm{n}$. Since the purpose of this appendix is to derive the magnon velocities, determined by parameters $\chi,c_{1},c_{2}$, and not the magnon gap which will depend sensitively on other (anisotropic) magnetic exchanges not included in Eq.(126), we henceforth drop the mass term in Eq.(G.4).
The last term is not rotationally invariant, therefore to recover the symmetry of Eq.(V.1) we select its rotation-invariant component. Technically this is simply done by writing $\bm{n}=R_{\theta}\tilde{\bm{n}}$ and $\bm{\partial}=R_{\theta}\tilde{\bm{\partial}}$, where $R_{\theta}$ is the 2D rotation matrix by an angle $\theta$ and $\tilde{\bm{n}},\tilde{\bm{\partial}}$ are the expression of ${\bm{n}},{\bm{\partial}}$ in a rotated orthogonal frame, coinciding with $(\hat{\mathbf{x}},\hat{\mathbf{y}})$ at $\theta=0$. We then average over the angle, $\frac{1}{2\pi}\int\text{d}\theta$. This yields

### v2

@@ Appendix E Solving Boltzmann’s equation and the displaced local equilibrium current
Here we derive the solution Eq.(35a) to the equation Eq.(III.3), to the first order in $\hbar$, equivalently in coordinate shifts $\delta\mathsf{X}$. To this order the “compensation term” $\mathsf{g}_{\rm d}^{\rm sj}(X,p)$ and out-of-equilibrium part $\mathsf{g}_{\rm d}^{\rm rta}(X,p)$ can be solved for independently, which we do following the steps of Ref.Sinitsyn et al. (2005) adapted to our generalized bosonic problem and notations.
@@ E.1 Displaced local equilibrium current
$\qquad\qquad\times\bigg[(\mathsf{g}_{\rm d})_{n,n}(X,p)-(\mathsf{g}_{\rm d})_{n^{\prime},n^{\prime}}(X,p^{\prime})-\delta\mathsf{X}^{\mu}_{n,n^{\prime}}(p,p^{\prime})\,\mathsf{\partial}_{X_{\mu}}T\;\partial_{T}f\!\left((\underline{\mathsf{K}}_{\rm d})_{n^{(}{}^{\prime}{}^{)},n^{(}{}^{\prime}{}^{)}}(X,p^{(}{}^{\prime}{}^{)})\right)\bigg]+\dots$
where $\dots$” contains terms higher-order in powers of $\hbar$ or beyond linear response, and the optional set of apostrophes in $\partial_{T}f$ follows from Eq.(102).
We now use the fact that Eq.(26), up to the approximations discussed in App.B, takes the form Eq.(31). Then, a solution to the collision integral is given by Eq.(35b) in the main text, which is clearly gauge invariant.
@@ E.2 Response in the relaxation time approximation (RTA)
$\qquad\qquad\times\left|\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})\right|^{2}\times\bigg[(\mathsf{g}_{\rm d}^{\rm rta})_{n,n}(X,p)-(\mathsf{g}_{\rm d}^{\rm rta})_{n^{\prime},n^{\prime}}(X,p^{\prime})\bigg].$
@@ E.3 Explicit writing of the side-jump current in the RTA
E.4 Generalizations
We now look for the equilibrium and out-of-equilibrium solutions $\mathsf{g}_{\rm d}^{\rm sj}$ and $\mathsf{g}_{\rm d}^{\rm rta}$ to the generalized kinetic equation, where we now assume explicitly only interband scattering ($n=n^{\prime}$ enforced):
[…]
$\quad+(\mathsf{g}_{\rm d})_{n,n}(p)-(\mathsf{g}_{\rm d})_{n,n}(p^{\prime})\bigg\}+\int_{p^{\prime}}\Upsilon^{\rm int}_{p,p^{\prime}}\Big[(\mathsf{g}_{\rm d})_{n,n}(p^{\prime})-(\mathsf{g}_{\rm d})_{n,n}(p)\Big],$ (109)
$(\mathsf{t}^{(\iota)}_{n,n})^{-1}(p,p^{\prime})$ $\equiv\frac{2\pi}{\hbar}\,\tilde{w}^{(\iota)}(p-p^{\prime})\;\delta\Big((\underline{\mathsf{K}}_{\rm d})_{n,n}(p^{\prime})-(\underline{\mathsf{K}}_{\rm d})_{n,n}(p)\Big)\;\big|\mathsf{M}^{(\iota)}_{n,n}(p,p^{\prime})\big|^{2}.$ (110)
First looking for the local equilibrium solution,where the lhs and the rhs vanish separately, one finds
$(\mathsf{g}_{\rm d}^{\rm sj})_{n,n}(p)$ $=-\Big(p_{\nu}\,\partial_{\mu}T\,\partial_{T}f\!\left((\underline{\mathsf{K}}_{\rm d})_{n,n}(p)\right)\;\sum_{(\iota)}(\mathsf{\Omega}^{\mathsf{W},(\iota)}_{p_{\mu}p_{\nu}})_{n,n}(p)\;\int_{p^{\prime}}(\mathsf{t}^{(\iota)}_{n,n})^{-1}(p,p^{\prime})\Big)\Big/\Big(\int_{p^{\prime}}\Big\{\Upsilon^{\rm int}_{p,p^{\prime}}+\sum_{(\iota)}(\mathsf{t}^{(\iota)}_{n,n})^{-1}(p,p^{\prime})\Big\}\Big).$ (111)
Then taking the order $O(\hbar^{0})$ on the rhs and restoring the gradient term on the lhs, one finds
$(\mathsf{g}_{\rm d}^{\rm rta})_{n,n}(p)$ $=\partial_{k_{\lambda}}\underline{\mathsf{K}}_{\rm d}(p)\;\partial_{\lambda}T\;\partial_{T}f(\underline{\mathsf{K}}_{\rm d}(p))\Big/\Big(\int_{p^{\prime}}\left(1-\hat{\rm p}\cdot\hat{\rm p}^{\prime}\right)\Big\{\Upsilon^{\rm int}_{p,p^{\prime}}+\sum_{(\iota)}(\mathsf{t}^{(\iota)}_{n,n})^{-1}(p,p^{\prime})\Big\}\Big).$ (112)
Assuming that $\Upsilon^{\rm int}_{p,p^{\prime}}$ just like $(\mathsf{t}^{(\iota)}_{n,n})^{-1}(p,p^{\prime})$ satisfies the usual assumptions of RTA (namely it only depends on the angle $\hat{\rm p}\cdot\hat{\rm p}^{\prime}$ and on the moduli $|p|=|p^{\prime}|$ constrained to be equal), one obtains Eq.(45) in the main text.
@@ F.1 Thermal Hall conductivity
Contrarily to the case of fermions in a random scalar potential, because $\mathsf{\Omega}^{\mathsf{W}}_{p_{\mu}p_{\nu}}$ is not antisymmetric a priori, these two currents are not identical a priori. The local current due to disorder, ${J}_{y}^{\rm dis}\equiv{J}_{y}^{\rm sj,rta}+{J}_{y}^{\rm gv,sj}$, is thus
@@ F.2 Electrical Hall effect
Appendix G Explicit formulae for numerical evaluation
Here we recapitulate how to evaluate Eqs.(1),(2) straightforwardly from a disordered bosonic theory of the form of Sec.V.1. Once the clean hamiltonian Eq.(47) and the disorder Hamiltonian Eq.(52) have been specified, Sec.V.1 provides explicit expressions for the dynamical matrix $\sf K$ and its disorder counterpart $\sf W$, both of which are $2N\times 2N$ matrices. Ultimately $\sf K$ and $\sf W$ fully determine the thermal Hall conductivities Eqs.(1),(2): we now detail the procedure to derive the latter.
The dynamical matrix is diagonalized as $\mathsf{S}_{0}^{-1}\mathsf{K}\mathsf{S}_{0}=\underline{\mathsf{K}}_{\rm d}$. Note that $\mathsf{S}_{0}$, whose columns are the bosonic eigenmodes, is generally not unitary (and is of course not uniquely defined). Note also that $\underline{\mathsf{K}}_{\rm d}$ is a diagonal matrix whose $2N$ entries are the $N$ bosonic energies (all positive) and their opposites (all negative). While Eqs.(1),(2) only involve summation over the $N$ physical bands, actual calculations of the curvatures $\mathsf{\Omega}_{p_{x}p_{y}},\mathsf{\Omega}^{\mathsf{W}}_{p_{x}p_{y}}$ can involve all $2N$ eigenvalues and eigenvectors.
For numerical evaluation, it is useful to resort to formulae for the Berry curvature ${\sf\Omega}_{p_{x}p_{y}}$ and its disorder counterpart $\tfrac{1}{2}(\mathsf{\Omega}^{\mathsf{W}}_{p_{x}p_{y}}-\mathsf{\Omega}^{\mathsf{W}}_{p_{y}p_{x}})$ that avoid computing momentum derivatives of gauge-dependent quantities. Such gauge-invariant formulae can be derived using the identity
$\left[\mathsf{S}_{0}^{-1}(\partial_{p_{x}}\mathsf{K})\mathsf{S}_{0}\right]_{n,n^{\prime}}=\left[\mathsf{S}_{0}^{-1}\partial_{p_{x}}\mathsf{S}_{0}\right]_{n,n^{\prime}}\left({[\underline{\mathsf{K}}_{\rm d}]_{n^{\prime},n^{\prime}}-[\underline{\mathsf{K}}_{\rm d}]_{n,n}}\right),$ (120)
which can be inverted for $n\neq n^{\prime}$. For instance it is well known that the Berry curvature can be recast into the form
$({\sf\Omega}_{p_{x}p_{y}})_{n,n}$ $={\rm Im}\left[\mathsf{S}_{0}^{-1}\partial_{p_{y}}\mathsf{S}_{0}\mathsf{S}_{0}^{-1}\partial_{p_{x}}\mathsf{S}_{0}-(x\leftrightarrow y)\right]_{n,n}$ (121)
$=\sum_{n^{\prime}\neq n}\frac{{\rm Im}\left(\left[\mathsf{S}_{0}^{-1}(\partial_{p_{x}}\mathsf{K})\mathsf{S}_{0}\right]_{n,n^{\prime}}\left[\mathsf{S}_{0}^{-1}(\partial_{p_{y}}\mathsf{K})\mathsf{S}_{0}\right]_{n^{\prime},n}\right)}{\left([\underline{\mathsf{K}}_{\rm d}]_{n^{\prime},n^{\prime}}-[\underline{\mathsf{K}}_{\rm d}]_{n,n}\right)^{2}}-(x\leftrightarrow y).$ (122)
Although only those indices $n$ corresponding to eigenmodes with positive eigenvalues are useful to compute Eq.(1), note that the summation $\sum_{n^{\prime}\neq n}$ involves all 2N12N-1 other indices.
A similar formula can be obtained for the antisymmetric part of ${\sf\Omega}^{\sf W}_{p_{x}p_{y}}$ (defined in Eq.(32)): one finds
$\tfrac{1}{2}(\mathsf{\Omega}^{\mathsf{W}}_{p_{x}p_{y}}-\mathsf{\Omega}^{\mathsf{W}}_{p_{y}p_{x}})_{n,n}=\tfrac{1}{2}({\sf\Omega}_{p_{x}p_{y}})_{n,n}$ (123)
$+\frac{1}{2}\sum_{n^{\prime}\neq n}\sum_{n^{\prime\prime}\neq n}{\rm Im}\bigg\{\Big(\left[\mathsf{S}_{0}^{-1}\partial_{p_{x}}\mathsf{S}_{0}\right]_{n^{\prime\prime},n}\left[\mathsf{S}_{0}^{-1}\partial_{p_{y}}\mathsf{S}_{0}\right]_{n,n^{\prime}}-(x\leftrightarrow y)\Big)\Big(\frac{\left[\mathsf{S}_{0}^{-1}\mathsf{W}\mathsf{S}_{0}\right]_{n^{\prime},n^{\prime\prime}}}{\left[\mathsf{S}_{0}^{-1}\mathsf{W}\mathsf{S}_{0}\right]_{n,n}}-\frac{\left[\mathsf{S}_{0}^{-1}\mathsf{W}\mathsf{S}_{0}\right]_{n^{\prime},n}\left[\mathsf{S}_{0}^{-1}\mathsf{W}\mathsf{S}_{0}\right]_{n,n^{\prime\prime}}}{\left[\mathsf{S}_{0}^{-1}\mathsf{W}\mathsf{S}_{0}\right]_{n,n}^{2}}\Big)\bigg\}$
$+\frac{1}{4}\sum_{n^{\prime}\neq n}{\rm Im}\bigg\{\frac{1}{\left[\mathsf{S}_{0}^{-1}\mathsf{W}\mathsf{S}_{0}\right]_{n,n}}\Big(\left[\mathsf{S}_{0}^{-1}\partial_{p_{y}}\mathsf{S}_{0}\right]_{n^{\prime},n}\left[\mathsf{S}_{0}^{-1}(\partial_{p_{x}}\mathsf{W})\mathsf{S}_{0}\right]_{n,n^{\prime}}+\left[\mathsf{S}_{0}^{-1}\partial_{p_{y}}\mathsf{S}_{0}\right]_{n,n^{\prime}}\left[\mathsf{S}_{0}^{-1}(\partial_{p_{x}}\mathsf{W})\mathsf{S}_{0}\right]_{n^{\prime},n}\Big)-(x\leftrightarrow y)\bigg\}$
$+\frac{1}{4}{\rm Im}\bigg\{\frac{\left[\mathsf{S}_{0}^{-1}(\partial_{p_{y}}\mathsf{W})\mathsf{S}_{0}\right]_{n,n}}{\left[\mathsf{S}_{0}^{-1}\mathsf{W}\mathsf{S}_{0}\right]_{n,n}^{2}}\sum_{n^{\prime}\neq n}\Big(\left[\mathsf{S}_{0}^{-1}\partial_{p_{x}}\mathsf{S}_{0}\right]_{n,n^{\prime}}\left[\mathsf{S}_{0}^{-1}\mathsf{W}\mathsf{S}_{0}\right]_{n^{\prime},n}+\left[\mathsf{S}_{0}^{-1}\partial_{p_{x}}\mathsf{S}_{0}\right]_{n^{\prime},n}\left[\mathsf{S}_{0}^{-1}\mathsf{W}\mathsf{S}_{0}\right]_{n,n^{\prime}}\Big)-(x\leftrightarrow y)\bigg\},$
which together with Eq.(120) provides a gauge-invariant formula suitable for numerical evaluation. This is the one we used to obtain our results in Sec.V.2. Again, although only those indices $n$ corresponding to eigenmodes with positive eigenvalues are useful to compute Eq.(2), note that the summations $\sum_{n^{\prime}\neq n}$ and $\sum_{n^{\prime\prime}\neq n}$ involve all 2N12N-1 other indices.
Appendix H Application to a low-energy bosonic theory – Details
H.1 Effective low-energy theory for magnons in an antiferromagnet
@@
In terms of $m_{i},n_{i}$ with $i=1,2$, and writing $\partial_{1}\equiv\partial_{x},\partial_{2}\equiv\partial_{y}$, the hamiltonian Eq.(V.3.1) reads
@@
Note – In the following, for notational clarity we write all the expressions in the $d_{3}=0$ and $\Delta=0$ case. While for numerical evaluations we indeed take $d_{3}=0$, they are still performed at finite $\Delta$ (with value given in Table 1).
H.2 Diagonalization of the clean problem
@@
Here we provide $\Psi_{1},\Psi_{2}$ to the first order in $\eta$ only. This is sufficient to compute the leading contribution to the Berry curvature $\mathsf{\Omega}_{p_{x}p_{y}}$ given above, Eq.(H.2), which is odd under time reversal, and thus odd in powers of $\eta$.
H.3 Explicit calculations in two given instances of disorder (b,c)
H.3.1 Case (b): a case with both intrinsic and extrinsic chirality
@@
$\mathsf{W}$ $={\sf\Gamma}\hat{W}=-\tilde{\eta}p^{2}\begin{bmatrix}\sigma^{y}&0\\ -i\eta p^{2}\text{\textbb{1}}&\sigma^{y}\end{bmatrix}$. (136)
@@
$\left[\mathsf{S}^{-1}(p)\begin{bmatrix}\sigma^{y}&0\\ -i\eta p^{2}\text{\textbb{1}}&\sigma^{y}\end{bmatrix}\mathsf{S}(p)\right]_{n,n}$ $=2\frac{\chi}{c_{2}}\eta\left(\varepsilon_{1}\;;\;-\varepsilon_{1}\;;\;\varepsilon_{2}\;;\;-\varepsilon_{2}\right)\;+O(\eta^{3}),$ (137)
@@
${\rm Im}\left[\mathsf{S}^{-1}(p)\begin{bmatrix}\sigma^{y}&0\\ -i\eta p^{2}\text{\textbb{1}}&\sigma^{y}\end{bmatrix}\partial_{p_{\mu}}\mathsf{S}(p)\right]_{n,n}$ $=\epsilon_{\mu,\overline{\mu}}(p_{\overline{\mu}}/p^{2})\;+O(\eta^{2}),$ (138)
@@
where $_{\mu}(p)$ and $^{\prime}_{\mu}(p)$ are real quantities which we do not write explicitly here.
H.3.2 Case (c): a case where the clean thermal Hall conductivity is zero
@@
$\mathsf{S}^{-1}(p)\mathsf{W}(p)\mathsf{S}(p)]_{n,n}=0,$
@@
H.4 Derivation from a microscopic spin model
@@
where $\mathtt{a}$} is the lattice parameter and $\dots$” contains terms that are higher order in gradients or beyond quadratic order in $\bm{m},\bm{n}$. Since the purpose of this appendix is to derive the magnon velocities, determined by parameters $\chi,c_{1},c_{2}$, and not the magnon gap which will depend sensitively on other (anisotropic) magnetic exchanges not included in Eq.(145), we henceforth drop the mass term in Eq.(H.4).
[…]